# Sideward contact tracing in an epidemic model with mixing groups

### v1

@@ 6 Numerical illustration
Throughout this section, we assume that the size of mixing event $C$ follows a geometric distribution, conditioned on being larger than 2 (following the choice in [ball2022epidemic]), that is,
$\mathbb{P}(C=c)=(1-\alpha)^{c-2}\alpha,\quad(c=2,3,...),$
with $\alpha=1/(\mu_{C}-1).$ In addition, we assume that the rate of natural recovery is $\gamma=1/7$.
First, we fix the average size of mixing events to be $\mu_{C}=5$. In Figure 6.1 we examine how the fraction of diagnosis $\delta/(\delta+\gamma)$ and tracing probability $p$ influence the effective reproduction number $R_{e}$. The infection probability $\pi_{c}$ is assumed first in Figure 1(a) independent of $c$, i.e. $\pi_{c}=0.5$ for all $c$, and then in Figure 1(b), $\pi_{c}={2}/{c}$. In each of the two cases, we fix $\beta$ such that the basic reproduction number $R_{0}$ in Equation (4) equals $3$. It is observed that $R_{e}$ is monotonically decreasing in both $\delta/(\delta+\gamma)$ and $p$ as expected. Additionally, $\delta/(\delta+\gamma)$ appears to have a stronger effect on reducing $R_{e}$ compared to $p$. This difference becomes more evident when $\pi_{c}={2}/{c}$. This is because, in comparison with the deterministic choice of $\pi_{c}=0.5$, fewer people, on average, become infected in a single mixing event of the same size and hence, more effort is required in sideward tracing to control the epidemic for a fixed reproduction number.
Further, we explore how the mean size of mixing events influences the effect of sideward tracing. We consider three situations of $\mu_{C}=5,10$ and $20$ while keeping $\delta=\gamma=1/7$ fixed. As shown in Figure 2(b), if we fix $R_{0}=3$ (different $\beta$ for different $\mu_{C}$), the reduction in the reproduction number due to contact tracing (i.e. $R_{0}-R_{e}$) is bigger for larger $\mu_{c}$ with the same tracing probability $p$. This is due to sideward tracing being more effective when more individuals get infected during the same event. This observation should not suggest that larger events are preferable, in fact, imposing limitations on the gathering size plays a bigger role in the reduction of the reproduction number. For instance, as shown in Figure 2(a), when we fix $\beta$ to be the same for all $\mu_{C}=5,10,20$, the case of $\mu_{C}=20,$ gives a much larger $R_{0}$ and even with perfect tracing ($p=1$), the reproduction number remains considerably high, well above 1.
@@
Figure 6.1: Heatmap of $R_{e}$ as function of $\delta/(\delta+\gamma)$ in $[0,0.99]$ and $p$ in $[0,1]$ with mean size of mixing event $\mu_{C}=5$, $\gamma=1/7$, in (a): $\pi_{c}=0.5$ for all $c$, in (b): $\pi_{c}=2/c$; and $\beta$ is chosen so that $R_{0}=3$.
@@
Figure 6.2: Plot of $R_{e}$ against the tracing probability $p$ for different mean size of events $\mu_{C}=5,10$ and $20$. The infection probability $\pi_{c}=2/c$ for all $c$, $\delta=\gamma=1/7$, and in (a), $\beta$ is fixed such that $R_{0}(\mu_{c}=5)=3$; in (b), $\beta$ is chosen specifically for each $\mu_{C}$ such that $R_{0}=3$.
@@ 7 Conclusion and discussion
A numerical illustration reveals that the fraction of diagnosis has a more significant impact on reducing $R_{e}$ than the tracing probability. This observation does not diminish the usefulness of sideward contact tracing. Note that increasing the fraction of diagnosis might be more challenging (especially in the beginning of the epidemic, requiring extensive testing efforts and more cost-effective tests) than simply increasing $p$ (e.g. during Covid-19 time in the UK, visitors/customers were encouraged to “check-in” by scanning QR code via NHS app [wymant2021epidemiological]). Furthermore, the impact of sideward tracing is more pronounced, resulting in a greater reduction in the reproductive number when the size of mixing events is larger. However, it remains crucial to impose limitations on gathering size. Without such restrictions, the reproductive number may remain excessively high, making it unfeasible to bring it below 1 even with perfect tracing.
@@
In this paper, we restrict our attention to sideward contact tracing only to analyse its effect separately from other types of contact tracing. One promising area for further investigation is combining sideward with conventional forward and backward tracing procedures which can be employed instead to trace infectors and infectees. It would be interesting to explore how these different tracing strategies can complement each other. For instance, in a scenario where an asymptomatic infector transmits the infection to two susceptibles during a mixing event, forward tracing may fail to identify the infectees as the infector remains undiagnosed. In such cases, sideward tracing becomes crucial if one of the infectees is diagnosed, enabling the identification of the remaining infectee. On the other hand, starting from one of the infectees, backward-forward tracing could potentially identify the asymptomatic infector but then fail to contact trace the other infectee before recovery. Integrating these tracing strategies is, however, analytically challenging, particularly due to the additional dependencies it introduces between the lifespans of siblings and their parents. Another possible extension of the present model making it more realistic is to incorporate delays between the diagnosis of the index case and the notification of other siblings. Additionally, incorporating heterogeneity into the model could provide deeper insights. For example, categorizing individuals based on their levels of social activity (high, normal, or low) could reveal more about the effectiveness of sideward tracing. More socially active individuals are more likely to attend events and consequently more likely to become infected and spread the infection to others. In this case, it would be important to trace the infector as well as the siblings.
@@
In conclusion, this paper highlights the potential effectiveness of sideward contact tracing and underscores that it can play a significant role in controlling epidemics, especially when relaxing the gathering size limitations.
@@ Acknowledgments
We would like to thank Tom Britton for valuable discussions and helpful comments throughout this work. Additionally, we extend our gratitude to Serik Sagitov for his constructive feedback. D.Z. acknowledges the Swedish Research Council (grant 2020-04744) for financial support. M.F. acknowledges the Knut and Alice Wallenberg Foundation (Program for Mathematics, grant 2020.072) for financial support.

### v2

@@ 6 Numerical results
    $C$ follows a geometric distribution, conditioned on being larger than 2, that is,
    $\mathbb{P}(C=c)=(1-\alpha)^{c-2}\alpha,\quad(c=2,3,...)$
    with $\alpha=\frac{1}{\mu_{C}-1}.$
  3. (iii)
    The event size is fixed: $C\equiv\mu_{C}$.
Through this examples, we investigate how the choice of mixing group size distribution and the forms of infection probability affects the effect of sideward contact tracing by examining its impact on the reduction of the reproduction number. The reproduction number $R_{e}$ is computed using its analytical expression in Equation (16). In our calculations, to ensure computational feasibility while maintaining approximation accuracy, the infinite sum over $c$ is truncated at $500$ in all the cases of $\mu_{C}$ and distributions of $C$.
Remark 6.1.
Whenever we compare scenarios corresponding to the two possible forms of $\pi_{c}$, i.e. $\pi_{c}\equiv\pi$ and $\pi_{c}={2}/{c}$, for a fair comparison, we choose the constant $\pi$ so that the average number of births per event is the same for both forms. More precisely, when $\pi_{c}={2}/{c}$, the average number of births per event is given by $\mathbb{E}[(\tilde{C}-1)\pi_{\tilde{C}}]$, implying the following condition for $\pi$:
$\frac{\pi\mathbb{E}[C(C-1)]}{\mu_{C}}=\frac{2(\mu_{C}-1)}{\mu_{C}}.$
Solving for $\pi$ leads to
$\pi=\frac{2(\mu_{C}-1)}{\mathbb{E}[C(C-1)]}.$
Specifically, if $C\sim Geo(\alpha),$ we have $\pi=\frac{1}{\mu_{C}-1}$; for $C\sim Log(\alpha)$, we get $\pi=\frac{2(\mu_{C}-1)\alpha}{\mu_{C}}$ and $C\equiv\mu_{C}$ gives $\pi=\frac{2}{\mu_{C}}.$
Remark 6.2.
Note that when $C$ is fixed, the varying infection probability $\pi_{c}=\frac{2}{c}$ if $c=\mu_{c};$ $\pi_{c}=0,$ otherwise. As a consequence, when the other parameters are the same, the effective reproduction number $R_{e}$ in terms of fixed infection probability $\pi=\frac{2}{\mu_{C}}$ is the same as the reproduction number $R_{e}$ for the varying $\pi_{c}$.
First, we plot the effective reproduction number $R_{e}$ as a function of the fraction of diagnosis, $\delta/(\delta+\gamma)$, and tracing probability $p$ for three distributions of $C$: logarithmic (Figure 1(a)), geometric (Figure 6.1) and being fixed (Figure 1(b)). In the left panels of each figure, the infection probability $\pi_{c}$ is assumed to be independent of $c$, i.e. $\pi_{c}\equiv\pi$; in the right panels, $\pi_{c}={2}/{c}$ for all $c$. The fixed infection probability $\pi$ is chosen specifically for each distribution of $C$ to result in the same average number of births per event as the case when $\pi_{c}={2}/{c}$ (see Remark 6.1). In either of the cases, we set $\gamma=1/7$ and choose $\beta$ such that the basic reproduction number $R_{0}$ in Equation (4) equals $3$.
As shown in Figure 1(a), 6.1 and 1(b), $R_{e}$ is monotonically decreasing in both $\delta/(\delta+\gamma)$ and $p$ as expected. Additionally, the diagnosis fraction $\delta/(\delta+\gamma)$ appears to have a stronger effect on reducing $R_{e}$ compared to tracing probability $p$ with this difference being slightly more pronounced when $\pi_{c}$ is fixed. The choice of group size distribution has a relatively weak effect on the influence of $\delta/(\delta+\gamma)$ and $p$ on reducing $R_{e}$, as the observed patterns are similar across distributions. Since the three figures (Figure 1(a), 6.1 and 1(b)) show similar behaviour, we report only one of them here and refer to Appendix B for the other two.
@@
Figure 6.1: Heatmaps of the reproduction number $R_{e}$ as function of $\delta/(\delta+\gamma)$ in $[0,0.99]$ and $p$ in $[0,1]$. The size of mixing event $C$ follows a geometric distribution with mean $\mu_{C}=20$, $\gamma=1/7$ and $\beta=3/266$. In the left panel: $\pi_{c}=2/c$ for $c\geq 2$, while $\pi_{c}\equiv 0.05$ on the right.
Next we focus on the reduction in the reproduction number, defined as $(R_{0}-R_{e})/R_{0}$. We note that this reduction is independent of rate $\beta$ based on the expressions in Equations (5) and (13). We further investigate how the reduction varies with the mean size of mixing events, $\mu_{C}$, the distributions of $C$ (geometric, logarithmic and fixed size) and the forms of $\pi_{C}$ ($2/c$ and fixed). Figure 6.2 illustrates the reduction for $\mu_{C}=5,10$ and $20$ while keeping $\delta=\gamma=1/7$ fixed. The fixed infection probability $\pi$ is selected for each distribution of $C$ with respective mean $\mu_{C}$ so that it gives the same average number of births in an event as for the corresponding $\pi_{c}={2}/{c}$ case. The results suggest that when $\pi_{c}={2}/{c}$, the distribution of $C$ does not significantly affect the reduction. In contrast, when the infection probability is fixed, the reduction becomes more sensitive to the choice of the distribution of $C$. Among the three distributions, logarithmic distribution results in the greatest reduction, followed by geometric distribution, with the fixed size showing the smallest reduction for the same mean size and tracing probability.
An explanation for the above observations is as follows. Recall that the expected number of infections per event, $\mathbb{E}[Y_{0}]$, is controlled to be the same across different distributions of $C$ with the same mean $\mu_{C}$. However, the variance $\mathrm{Var}[Y_{0}]$ depends on the distribution of $C$. A higher variance in $Y_{0}$ means that some events have higher probability of resulting in a relatively large number of infections (“super-spreading events”). These are precisely the events which are best targeted by sideward contract tracing. When the infection probability $\pi_{c}\equiv\pi$ is fixed, $\mathrm{Var}[Y_{0}]$ increases with the variance of $C$, which is higher under the logarithmic distribution, lower under the geometric and equal to zero in the deterministic case, explaining the different effects of sideward tracing under the three considered distributions. On the other hand, when the infection probability takes the specific form $\pi_{c}=2/c$, a “balancing” effect is introduced. This reduces the dependence of $\mathrm{Var}[Y_{0}]$ on the distribution of $C$, making the impact of sideward tracing more uniform across different distributions of $C$.
Figure 6.2: Plot of the relative reduction $r$ as a function of the tracing probability $p$ for different mean event size $\mu_{C}=5,10$ and $20$. We fix $\delta=\gamma=1/7$ and in the upper panel, the infection probability $\pi_{c}={2}/{c}$ for all $c$; in the lower panel, $\pi_{c}\equiv\pi$ is chosen specifically for each distribution of $C$ to result in the same average number of infections per event as the $\pi_{c}={2}/{c}$ case.
Crucially, we observe that for all mixing event distributions and for all forms of $\pi_{C}$, a larger $\mu_{c}$ consistently results in a greater reduction corresponding to a certain tracing probability $p$. This is due to sideward tracing being more effective when more individuals get infected during the same event. Nevertheless, this observation should not suggest that larger events are preferable, in fact, imposing limitations on the gathering size plays a bigger role in the reduction of the reproduction number. As shown in Figure 6.3, if $\beta$ is fixed across all scenarios ($\mu_{C}=5,10$ and $20$), the case of $\mu_{C}=20$ gives a much larger $R_{0}$. In this case, even under perfect tracing ($p=1$), when the reduction is substantial, the reproduction number $R_{e}$ remains considerably high, well above 1.
@@
Figure 6.3: Plot of the reproduction number $R_{e}$ as a function of the tracing probability $p$ for different mean event size $\mu_{C}=5,10$ and $20$. We fix $\delta=\gamma=1/7$ and $\beta=3/28$ such that $R_{0}=3$ with $\mu_{C}=5$. In the upper panel, the infection probability $\pi_{c}={2}/{c}$ for all $c$; in the lower panel, $\pi_{c}\equiv\pi$ is chosen specifically for each distribution of $C$ to result in the same average number of infections per event as the $\pi_{c}={2}/{c}$ case.
@@ 7 Conclusion and discussion
Using the obtained theoretical result, we performed a numerical investigation to show variations in the effect of sideward contact tracing depending on various quantities. Our numerical results reveal that the fraction of diagnosis has a greater impact on reducing $R_{e}$ compared to the tracing probability $p$, in absolute terms. However, it should be noted that these two quantities are intertwined and hardly directly comparable. The effectiveness of tracing is inherently dependent on successful diagnosis, as tracing cannot occur if cases are not diagnosed. Increasing the fraction of diagnosed cases presents significant challenges, particularly in the beginning of an epidemic, as it requires extensive testing efforts and the development of cost-effective tests. Similarly, increasing the fraction of traced contacts is also challenging. As studied by [11], the system can become overwhelmed when large numbers of contacts need to be traced rapidly, especially during periods of high case importation.
Furthermore, our numerical results also show that the impact of sideward tracing is more pronounced, resulting in a greater reduction in the reproductive number, when the size of mixing events is larger. This emphasizes the potential of sideward tracing as a control measure in connection with large gatherings. We also highlight the limitations of sideward tracing: without restrictions on the gathering size, the reproductive number may remain excessively high, making it unfeasible to bring it below 1 even with perfect tracing.
The theoretical results in this paper also provide technical tools for further analyses. For example, deriving the growth rate of the macro process would allow to characterize the growth rate of the epidemic as well as the generation time distribution in terms of macro individuals, see Appendix A for more details. Previous research [8] has shown that some preventive measures (as isolation, mass testing, forward and backward contact tracing) affect the generation time distribution leading to biased estimates of the reproduction number. This suggests that a similar study on the effect of sideward contact tracing on the generation time distribution and resulting biases might yield interesting results.
@@
In this paper, we restricted our attention to sideward contact tracing in order to analyse its effect separately from other types of contact tracing. One promising area for further investigation arises from combining sideward with conventional forward and backward tracing procedures which can be employed instead to trace infectors and infectees. It would be interesting to explore how these three different tracing strategies can complement each other in the framework of this paper, similarly to [15] who investigated the effectiveness of combining tracing mechanisms in the framework of temporal networks. For instance, in a scenario where an asymptomatic infector transmits the infection to two susceptibles during a mixing event, forward tracing may fail to identify the infectees as the infector remains undiagnosed. In such cases, sideward tracing becomes crucial if one of the infectees is diagnosed, enabling the identification of the remaining infectee. On the other hand, starting from one of the infectees, backward-forward tracing could potentially identify the asymptomatic infector but then fail to contact trace the other infectee before recovery. Integrating these tracing strategies in our branching-processes framework is analytically challenging, particularly due to the additional dependencies it introduces between the lifespans of siblings and their parents.
Another possible extension of our tracing model is to incorporate delays between the diagnosis of the index case and the notification of other siblings. The impact of such tracing delays has been extensively studied in traditional contact tracing frameworks. For instance, [4] analysed the impact of delays in forward tracing, while [16] investigated the effects of delays in both forward and backward tracing. Furthermore, [13] demonstrated that minimizing tracing delays and optimizing tracing coverage are crucial for the success of a contact tracing strategy.
Additionally, incorporating heterogeneity into the model could provide deeper insights. For example, [12] highlights the effectiveness of backward contact tracing, particularly in leveraging the heterogeneity of network structures. One interesting extension to our present model can be categorizing individuals based on their levels of social activity (high, normal, or low) could reveal more about the effectiveness of sideward tracing. More socially active individuals are more likely to attend events and consequently more likely to become infected and spread the infection to others. In this case, it would be important to trace the infector as well as the siblings.
@@
In conclusion, this paper highlights the potential effectiveness of sideward contact tracing. We show that it can play a significant role in controlling epidemics, especially when relaxing the gathering size limitations, and we underline its limitations.
@@ Acknowledgments
We would like to thank Tom Britton for valuable discussions and helpful comments throughout this work. Additionally, we extend our gratitude to Serik Sagitov for his constructive feedback. Most of the work was conducted while D.Z. was pursuing her PhD at the Department of Mathematics, Stockholm University, 106 91 Stockholm, Sweden.
Funding
D.Z. acknowledges the Swedish Research Council (grant 2020-04744) for financial support. M.F. acknowledges the Knut and Alice Wallenberg Foundation (Program for Mathematics, grant 2020.072) for financial support.
Appendix A Growth rate of the macro process $\mathcal{M}$
The macro-process $\mathcal{M}$ constitutes a technical framework that can be used for further analyses. We report here some additional technical information which we deem useful.
Being $\lambda(t)$ the birth rate of the macro branching process, it is possible to easily express the corresponding generation time distribution by normalising $\mathbb{E}\left[\lambda(t)\right]$ so that it integrates to $1$ (dividing by the reproduction number). Furthermore, the Malthusian parameter, or growth rate, $r_{\mathcal{M}}$ can be expressed as the unique solution of
[…]
where $\sigma_{i}(s)$ is the time of birth of the $i^{th}$ macro-individual alive at time $s$, with $s-\sigma_{i}(s)$ thus being its age. By interpreting the size process $Y$ as a “characteristic” of a macro-individual, the expression above allows us to interpret $I_{\mathcal{B}_{CT}}(s)$ as the “total characteristic” of $\mathcal{M}$ at time $s$, as in [10]. Then, the classical theory of Crump-Mode-Jagers branching processes [10] implies in the supercritical case that, for large times, $I_{\mathcal{M}}(s)\approx e^{r_{\mathcal{M}}s}W$, with $W$ being a positive random variable; and that $I_{\mathcal{B}_{CT}}(s)\approx e^{r_{\mathcal{M}}s}m_{Y}W$, with $m_{Y}$ being a positive constant.
Appendix B Additional numerical results
Figure B.1: Heatmaps of the reproduction number $R_{e}$ as function of $\delta/(\delta+\gamma)$ in $[0,0.99]$ and $p$ in $[0,1]$, where $\gamma=1/7$ and $\beta=3/266$. The size of mixing event $C$ follows a logarithmic distribution with mean $\mu_{C}=20$ in 1(a), where in the left panel: $\pi_{c}=2/c$ for $c\geq 2$ and $\pi_{c}\equiv 0.03$ on the right. In 1(b), the event size is fixed: $C\equiv 20$ and hence $\pi_{c}\equiv\pi=0.1$.